40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2348     Entry time: Mon May 20 02:23:14 2019
Author: Anjali 
Type: Update 
Category: 2micronLasers 
Subject: Frequency noise measurement of 2 micron source using PLL 
  • Attachment # 1 shows the schematic of the experimental setup for the frequency noise measurement of 2-micron laser source using PLL. Instead of Brimrose driver, another Marconi is used to provide the RF power to the AOM. We know from the characterisation of AOM that we need to give RF power of 28 dBm at 78.8 MHz to achieve maximum diffraction efficiency to the first order port of AOM. The maximum output power from Marconi is 13 dBm. Hence, we used another RF amplifier (ZHL-3-A+) to amplify the RF power from Marconi. We initially tested the RF output from RF amplifier on spectrum analyser (RF power fed into spectrum analyser with proper attenuation in the path) and adjusted the RF frequency and power in Marconi such that we get 28 dBm output power from the RF amplifier at 78.8 MHz. The two marconis are set such that they are share the same time standard.

  • Now, the output power from the photodetector in MZI (Laser diode operated at input current of 90 mA) is fed into the RF input port of the mixer, instead of AFG. The 600 Ohm output of SR 560 is observed on oscilloscope and SR 785 simultaneously.

  • We observed dc line in the oscilloscope when the gain in SR 560 is set to 13 dB (20 times). Gain value below this ( 10 dB) or above this (17 dB) was showing oscillations in the oscilloscope with frequency varying with the actuation slope in Marconi. Attachment #2 shows the frequency noise measurement from SR 785 (V/rt Hz value from SR 785 multiplied with the actuation slope).  

  • It is observed that, the time domain trace on the oscilloscope was not very stable. In between, we could see the oscillation was popping up. Also, the trace on SR 785 was swinging a lot (attached the video). As we observed in the case of AFG, the FM noise measured increses with the value of actuation slope in Marconi. 

  • In this case, the RF power that is fed into the RF port of the mixer is very small (~ -40 dBm) compared to our previous experiment with AFG. So, I would like to repeat the sample experiment (locking AFG to Marconi) with AFG set to RF power comparable with that from the actual experiment.I should then find out the unity gain frequency of that particular combination of gain and actuation slope,which would help us to find the frequency range upto which the PLL measurment is valid.

  •  I also need to measure the actual delay line length. I will also clean up the fiber connectors again and we can also use the FC/UPC to FC/APC patch cord for the detector after it is arrived. I still must understand the results better.Since we are using Non-PM fibers, the polarisation fluctuation might have also affected the measurement .Kindly give me further suggestions.

Attachment 1: FM_noise_measurement_setup.png  126 kB  | Hide | Hide all
FM_noise_measurement_setup.png
Attachment 2: 2_micron_FM_noise.pdf  97 kB  | Hide | Hide all
2_micron_FM_noise.pdf
Attachment 3: Oscilloscope.mp4  6.960 MB  Uploaded Mon May 20 11:48:52 2019
Attachment 4: SR_785.mp4  9.002 MB  Uploaded Mon May 20 11:49:12 2019
ELOG V3.1.3-